# MokA: Multimodal Low-Rank Adaptation for MLLMs

### v1

@@ 1 Introduction
Large language models (LLMs) have gained remarkable popularity due to their impressive ability to understand and generate content. To extend their capabilities to more general multimodal scenarios, recent advancements of Multimodal Large Language Models (MLLMs)Β [35, 37, 18] have focused on aligning other modalities, such as images, with text tokens, thereby equipping LLMs with the ability to interpret and process content of other modalities. However, due to the massive parameter scale of LLMs, fully fine-tuning such models on downstream tasks is computationally prohibitive and inefficient in most cases.
@@
A promising direction has emerged in the field of LLM fine-tuning before, which involves selectively updating a subset of parameters rather than the full model. These Parameter-Efficient Fine-Tuning (PEFT) strategies have seen widespread adoption and have been successfully extended to the fine-tuning of MLLMs. In particular, LoRAΒ [11] and its variants, which assume that over-parameterized models in fact reside on a low intrinsic dimension, have been broadly appliedΒ [5, 6, 36], demonstrating strong adaptability and efficiency. However, the development of efficient multimodal LLM fine-tuning is at present obscured by a βdark cloudβ: most current methods are directly borrowed from LLMs, often overlooking the fundamental differences of multimodal scenarios. Indeed, prior studies in multimodal learning have demonstrated that the inherent heterogeneity of different modalities necessitates modality-specific utilization strategies, rather than a fully unified wayΒ [22, 31].
@@
To this end, we are motivated to observe the fine-tuning efficacy of the widely used LoRA strategy. A common MLLM fine-tuning framework is shown inΒ 1(a): encoded representation of non-text modality (e.g., audio or visual) is first aligned with the text embedding space via a projector (usually Q-former or MLP), after which the resulting multimodal tokens are integrated and processed jointly by the LLM. In the efficient fine-tuning case, the LLM backbone is frozen, and parameters of additional LoRA modules are optimized. 1(b) provides the sketch of classic LoRA module. and matrices are shared across different modalities.
@@
To further observe how well tokens of different modalities are utilized, we conduct partial modality inference experiments. The training stage retains the original setting, wherein all multimodal tokens are processed by the LoRA module. Specifically, we evaluate the modelβs performance when only tokens from a selected modality are passed through the LoRA adaptation pathway at the first generation step during inference. As illustrated inΒ 1(c), visual token inference is used as an example. And it should be noted that the pre-trained weights still receive full tokens of all modalities. Results shown inΒ FigureΒ 1d-1f demonstrate a surprising phenomenon across three representative multimodal scenarios, audio-visual-text case, visual-text case, and speech-text case. Text token inference can achieve quite comparable performance to the regular full modalities case. However, Non-text token inference (e.g., audio or visual) leads to a noticeable drop in performance.
@@ 2.1 Rethinking of low-rank adaptation in the multimodal scenario
LoRAΒ [11] is based on the assumption that the weight updates during fine-tuning lie in a subspace of low βintrinsic rank.β Rather than updating the entire pre-trained weight matrix directly, LoRA introduces a low-rank decomposition approach, where the update to a pre-trained matrix is parameterized as the product of two much smaller matrices: and , with . The resulting fine-tuned weight matrix is given by . Therefore, for , the modified update forward pass yields:
Here, remains fixed during training, while only the matrices and are learned. To ensure stable training, is initialized using a uniform Kaiming distributionΒ [10], and is initialized to zero, leading to an initial update at the beginning of fine-tuning. LoRAΒ [11] and its variants have been extensively employed in the parameter-efficient fine-tuning of MLLMsΒ [5, 6, 36]. These methods typically employ shared parameters to uniformly process tokens from all modalities, implicitly facilitating cross-modal interactions during adaptation. However, our empirical results reveal that such shared tuning leads to limited utilization of all modalities. This highlights the need to consider individual modality during fine-tuning.
@@ 2.2 Multimodal low-rank Adaptation (MokA)
Based on the above perspective, we propose Multimodal low-rank Adaptation (MokA) strategy, a parameter-efficient fine-tuning method tailored for the multimodal nature of MLLMs. Considering the efficiency advantage of LoRA, MokA retains the core idea of low-rank adaptation, but redefines the roles of the projection matrices and to better reflect the characteristics of multimodal scenarios. By unimodal compression and explicitly reinforcing cross-modal interaction, MokA enables both unimodal and cross-modal adaptation, leading to more effective fine-tuning of MLLMs. The overall structure of MokA is depicted inΒ FigureΒ 2.
@@ 2.2.2 Task-centric cross-attention
In this case, <audio> and <visual> provide the event information. βPlease answer the question: which clarinet makes the sound first?β describes the concrete task for LLMs. Successfully answering such questions relies on effectively capturing the semantic association between the task description conveyed by text tokens and the event cues provided by non-text tokens. Therefore, it becomes intuitive and necessary to explicitly emphasize the most relevant cross-modal information to support accurate reasoning. Since unimodal information has been extracted individually after the processing of unimodal matrices , this stage is well-suited for introducing cross-modal interaction. Additionally, as the token embeddings are projected into a low-rank space, the computational burden of performing cross-modal interaction is significantly reduced. Hence, we place the cross-attention part after the low-rank compression to ensure both effectiveness and efficiency. The concrete attention mechanism is illustrated inΒ FigureΒ 3, and is conducted as follows:
@@
Finally, the sequence after cross-attention is:
Here we use to simply denote the cross-attention between modality and text. It should be noted that while we adopt a cross-attention module to explicitly enhance the interaction between text and non-text tokens, alternative designs that serve a similar purpose can also be considered. Further discussion is provided inΒ SectionΒ 4.4. In addition, in MokA, linear projections (, , and ) are not included in the cross-attention module, since low-rank matrices of each modality actually can be considered as the linear projection in attention in this case. We also provide more discussion and comparison inΒ AppendixΒ B.
@@ 2.2.4 Overview
where denotes the zero vector of dimension , since text-token remains unchanged after cross-attention. During fine-tuning, remains unchanged, with and being subject to optimization. Also, is initialized using the uniform Kaiming distributionΒ [10], while is initialized to zero. It leads to an initial update at the beginning of fine-tuning, to provide a smooth starting point.
Based onΒ EquationΒ 13, MokA ensures both unimodal and cross-modal adaptation, offering a more tailored solution for fine-tuning MLLMs.
@@ 3.1 Implement details
Our framework follows the common MLLM framework as illustrated inΒ 1(a), but with MokA strategy. Text input is processed by the corresponding LLM tokenizer, and non-text input is first encoded by its encoder, and then aligned with the text embedding space via a projector. Here we use Q-former followed by a two-layer MLP as the projector. Finally, all tokens are fed into LLM.
For the visual branch of audio-visual-text and visual-text scenarios, we use CLIP-ViT/L-14Β [23] as the visual encoder to extract the last layer patch level embedding of each frame or image. For the audio branch of the audio-visual-text scenario, we use the BEATsΒ [4] encoder to extract features. For the speech branch of the speech-text scenario, OpenAIβs Whisper modelΒ [24] is used. The number of query tokens in Q-Former of all branches is .
@@ 3.2 Training procedure and benchmarks
Pre-training: LLM backbone is frozen. Projectors are trainable for cross-modal alignment. For the visual branch of audio-visual-text and visual-text scenarios, trainable modules are trained on video-LLaVAΒ [17] dataset, including the video captioning and the image captioning tasks. For the audio branch of the audio-visual-text scenario, trainable modules are trained on AudioCapsΒ [12] dataset on the audio captioning task. For the speech branch of the speech-text scenario, trainable modules are trained on GigaSpeech-MΒ [3] dataset on the speech recognition task. During pre-training, each branch is trained for one epoch, using the AdamW optimizer with a cosine learning rate schedule. The initial learning rate is with a warmup ratio of .
Instruction-tuning: At this stage, we train the model on downstream tasks in different scenarios. Trainable parameters include all projectors and our MokA module. For the audio-visual-text case, the model is fine-tuned on the train set of MUSIC-AVQAΒ [15], and AVEΒ [28], respectively. For the visual-text case, the model is fine-tuned on the LLaVA-Instruct-150KΒ [18]. For the speech-text case, the model is fine-tuned on the LibriSpeechΒ [21], using the annotations provided byΒ [26]. The visual-text case is trained for one epoch, and other cases are trained for three epochs. Rank of low-rank matrices is . The remaining settings are the same as the first stage.
@@
Β Β Β Audio-visual-text: in-domain test set of MUSIC-AVQA and AVE dataset.
Β Β Β Visual-text: public benchmarks: MMEpercepΒ [8], MMBenchΒ [20], POPEΒ [16], SEED-BenchΒ [14].
Β Β Β Speech-text: public benchmarks: MMAUmin-speechΒ [25] as well as the foundation subset of
Β Β Β Β Β Β AIR-Benchspeech-enΒ [34].
Large Language Model. For all three cases, LLaMA-2-7b-ChatΒ [29], LLaMA-3-8B-InstructΒ [9], and Qwen2-7B-InstructΒ [33] are used as the LLM base model, respectively. For the audio-visual-text case, Qwen2.5-VL-7B-InstructΒ [2] is also used as the LLM base model. Throughout the training process, weights of LLM are kept frozen. More experiments of Qwen3 are provided inΒ AppendixΒ A.
@@ 4.1 Audio-visual-text scenario
| LoRAMoEΒ [7] | 73.57 | 72.81 | |||
| DoRAΒ [19] | 73.97 | 72.18 | 1 | 1 | |
| HydraLoRAΒ [27] | 74.12 | 72.27 | 1 | ||
| Uni-modal LoRAΒ [1] | 74.37 | 71.44 | |||
@@
| LoRAMoEΒ [7] | 73.48 | 72.83 | |||
| DoRAΒ [19] | 73.29 | 72.91 | 1 | 1 | |
| HydraLoRAΒ [27] | 73.14 | 72.59 | 1 | ||
| Uni-modal LoRAΒ [1] | 73.62 | 73.14 | |||
@@
| LoRAMoEΒ [7] | 73.28 | 71.91 | |||
| DoRAΒ [19] | 73.37 | 71.06 | 1 | 1 | |
| HydraLoRAΒ [27] | 73.04 | 71.26 | 1 | ||
| Uni-modal LoRAΒ [1] | 73.46 | 72.11 | |||
@@
To validate the effectiveness of our MokA fine-tuning strategy, we compare it with LoRAΒ [11] and its variants, including multiple LoRA, LoRAMoEΒ [7], DoRAΒ [19], HydraLoRAΒ [27], Uni-modal LoRAΒ [1]. In addition, we also compare with two additional baselines, whose frameworks are provided inΒ FigureΒ 4. Concretely, the Uni LoRA + MM LoRA strategy employs unimodal low-rank matrices to extract unimodal information independently, while incorporating an additional fully shared multimodal LoRA module to implicitly promote cross-modal interaction. The Uni LoRA + MM LoRA + Gate variant further introduces a gating mechanism to dynamically integrate the outputs of the Uni LoRA and MM LoRA branches for improved fusion. These two baselines incorporate our multimodal-aware basis that ensures both unimodal and cross-modal adaptation, but involve more parameters and offer limited cross-modal interaction. Based on the results inΒ TableΒ 1, we can have the following observations:
Our proposed MokA method achieves the superior overall performance across multiple audio-visual-text datasets, consistently outperforming other baselines and compared methods. While MokA introduces a slight increase in parameter scale compared to standard LoRA, this does not account for the observed performance improvements. Based on theΒ TableΒ 1, multiple LoRA, a baseline that uses A matrices and A matrices, underperforms both standard LoRA and MokA. Simply increasing the number of low-rank matrices does not necessarily lead to better fine-tuning performance. This suggests that MokAβs advantage stems not from parameter quantity, but from its insurance for both unimodal and multimodal adaptation.
@@
| LoRAMoEΒ [7] | 938.52 | 51.98 | 71.15 | 39.13 | |||
| DoRAΒ [19] | 786.47 | 51.31 | 71.07 | 38.96 | 1 | 1 | |
| HydraLoRAΒ [27] | 774.47 | 47.33 | 70.87 | 38.81 | 1 | ||
| Uni-modal LoRAΒ [1] | 992.31 | 51.98 | 72.24 | 39.27 | |||
@@
| LoRAMoEΒ [7] | 1157.39 | 57.29 | 81.29 | 56.39 | N | N | |
| DoRAΒ [19] | 1024.42 | 56.19 | 80.75 | 55.03 | 1 | 1 | |
| HydraLoRAΒ [27] | 1098.25 | 56.42 | 81.34 | 54.67 | 1 | ||
| Uni-modal LoRAΒ [1] | 1189.47 | 57.39 | 81.12 | 56.21 | |||

### v2

@@ 1 Introduction
Large language models (LLMs) have gained remarkable popularity due to their impressive ability to understand and generate content. To extend their capabilities to more general multimodal scenarios, recent advancements of Multimodal Large Language Models (MLLMs) [ye2023mplug, zhu2023minigpt, liu2023visual] have focused on aligning other modalities, such as images, with text tokens, thereby equipping LLMs with the ability to interpret and process content of other modalities. However, due to the massive parameter scale of LLMs, fully fine-tuning such models on downstream tasks is computationally prohibitive and inefficient in most cases.
@@
A promising direction has emerged in the field of LLM fine-tuning before, which involves selectively updating a subset of parameters rather than the full model. These Parameter-Efficient Fine-Tuning (PEFT) strategies have seen widespread adoption and have been successfully extended to the fine-tuning of MLLMs. In particular, LoRA [hu2022lora] and its variants, which assume that over-parameterized models in fact reside on a low intrinsic dimension, have been broadly applied [chen2024llava, chen2024salm, yeo2024visual], demonstrating strong adaptability and efficiency. However, the development of efficient multimodal LLM fine-tuning is at present obscured by a “dark cloud”: most current methods are directly borrowed from LLMs, often overlooking the fundamental differences of multimodal scenarios. Indeed, prior studies in multimodal learning have demonstrated that the inherent heterogeneity of different modalities necessitates modality-specific utilization strategies, rather than a fully unified way [peng2022balanced, wei2024fly].
@@
To this end, we are motivated to observe the fine-tuning efficacy of the widely used LoRA strategy. A common MLLM fine-tuning framework is shown in 1(a): encoded representation of non-text modality (e.g., audio or visual) is first aligned with the text embedding space via a projector (usually Q-former or MLP), after which the resulting multimodal tokens are integrated and processed jointly by the LLM. In the efficient fine-tuning case, the LLM backbone is frozen, and parameters of additional LoRA modules are optimized. 1(b) provides the sketch of classic LoRA module. and matrices are shared across different modalities.
@@
To further observe how well tokens of different modalities are utilized, we conduct partial modality inference experiments. The training stage retains the original setting, wherein all multimodal tokens are processed by the LoRA module. Specifically, we evaluate the model’s performance when only tokens from a selected modality are passed through the LoRA adaptation pathway at the prefilling stage during inference. As illustrated in 1(c), visual token inference is used as an example. And it should be noted that the pre-trained weights still receive full tokens of all modalities. Results shown in Figure 1d-1f demonstrate a surprising phenomenon across three representative multimodal scenarios, audio-visual-text case, visual-text case, and speech-text case. Text token inference can achieve quite comparable performance to the regular full modalities case. However, Non-text token inference (e.g., audio or visual) leads to a noticeable drop in performance.
@@ 2.1 Rethinking of low-rank adaptation in the multimodal scenario
LoRA [hu2022lora] is based on the assumption that the weight updates during fine-tuning lie in a subspace of low “intrinsic rank.” Rather than updating the entire pre-trained weight matrix directly, LoRA introduces a low-rank decomposition approach, where the update to a pre-trained matrix is parameterized as the product of two much smaller matrices: and , with . The resulting fine-tuned weight matrix is given by . Therefore, for , the modified update forward pass yields:
Here, remains fixed during training, while only the matrices and are learned. To ensure stable training, is initialized using a uniform Kaiming distribution [he2015delving], and is initialized to zero, leading to an initial update at the beginning of fine-tuning. LoRA [hu2022lora] and its variants have been extensively employed in the parameter-efficient fine-tuning of MLLMs [chen2024llava, chen2024salm, yeo2024visual]. These methods typically employ shared parameters to uniformly process tokens from all modalities, implicitly facilitating cross-modal interactions during adaptation. However, our empirical results reveal that such shared tuning leads to limited utilization of all modalities. This highlights the need to consider individual modality during fine-tuning.
@@ 2.2 Multimodal low-rank Adaptation (MokA)
Based on the above perspective, we propose Multimodal low-rank Adaptation (MokA) strategy, a parameter-efficient fine-tuning method tailored for the multimodal nature of MLLMs. Considering the efficiency advantage of LoRA, MokA retains the core idea of low-rank adaptation, but redefines the roles of the projection matrices and to better reflect the characteristics of multimodal scenarios. By unimodal compression and explicitly reinforcing cross-modal interaction, MokA enables both unimodal and cross-modal adaptation, leading to more effective fine-tuning of MLLMs. The overall structure of MokA is depicted in Figure 2.
@@ 2.2.2 Task-centric cross-attention
In this case, <audio> and <visual> provide the event information. “Please answer the question: which clarinet makes the sound first?” describes the concrete task for LLMs. Successfully answering such questions relies on effectively capturing the semantic association between the task description conveyed by text tokens and the event cues provided by non-text tokens. Therefore, it becomes intuitive and necessary to explicitly emphasize the most relevant cross-modal information to support accurate reasoning. Since unimodal information has been extracted individually after the processing of unimodal matrices , this stage is well-suited for introducing cross-modal interaction. Additionally, as the token embeddings are projected into a low-rank space, the computational burden of performing cross-modal interaction is significantly reduced. Hence, we place the cross-attention part after the low-rank compression to ensure both effectiveness and efficiency. The concrete attention mechanism is illustrated in Figure 3, and is conducted as follows:
@@
where and are the hyperparameters that control the strength of explicit cross-modal interaction. Finally, the sequence after cross-attention is:
Here we use to simply denote the cross-attention between modality and text. It should be noted that while we adopt a cross-attention module to explicitly enhance the interaction between text and non-text tokens, alternative designs that serve a similar purpose can also be considered. Further discussion is provided in Section 4.4. In addition, in MokA, linear projections (, , and ) are not included in the cross-attention module, since low-rank matrices of each modality actually can be considered as the linear projection in attention in this case. We also provide more discussion and comparison in Appendix B.
@@ 2.2.4 Overview
where denotes the zero vector of dimension , since text-token remains unchanged after cross-attention. During fine-tuning, remains unchanged, with and being subject to optimization. Also, is initialized using the uniform Kaiming distribution [he2015delving], while is initialized to zero. It leads to an initial update at the beginning of fine-tuning, to provide a smooth starting point.
Based on Equation 13, MokA ensures both unimodal and cross-modal adaptation, offering a more tailored solution for fine-tuning MLLMs.
@@ 3.1 Implement details
Our framework follows the common MLLM framework as illustrated in 1(a), but with MokA strategy. Text input is processed by the corresponding LLM tokenizer, and non-text input is first encoded by its encoder, and then aligned with the text embedding space via a projector. Here we use Q-former followed by a two-layer MLP as the projector. Finally, all tokens are fed into LLM.
For the visual branch of audio-visual-text and visual-text scenarios, we use CLIP-ViT/L-14 [radford2021learning] as the visual encoder to extract the last layer patch level embedding of each frame or image. For the audio branch of the audio-visual-text scenario, we use the BEATs [chen2022beats] encoder to extract features. For the speech branch of the speech-text scenario, OpenAI’s Whisper model [radford2023robust] is used. The number of query tokens in Q-Former of all branches is .
@@ 3.2 Training procedure and benchmarks
Pre-training: LLM backbone is frozen. Projectors are trainable for cross-modal alignment. For the visual branch of audio-visual-text and visual-text scenarios, trainable modules are trained on video-LLaVA [lin2023video] dataset, including the video captioning and the image captioning tasks. For the audio branch of the audio-visual-text scenario, trainable modules are trained on AudioCaps [kim2019audiocaps] dataset on the audio captioning task. For the speech branch of the speech-text scenario, trainable modules are trained on GigaSpeech-M [chen2021gigaspeech] dataset on the speech recognition task. During pre-training, each branch is trained for one epoch, using the AdamW optimizer with a cosine learning rate schedule. The initial learning rate is with a warmup ratio of .
Instruction-tuning: At this stage, we train the model on downstream tasks in different scenarios. Trainable parameters include all projectors and our MokA module. For the audio-visual-text case, the model is fine-tuned on the train set of MUSIC-AVQA [li2022learning], and AVE [tian2018audio], respectively. For the visual-text case, the model is fine-tuned on the LLaVA-Instruct-150K [liu2023visual]. For the speech-text case, the model is fine-tuned on the LibriSpeech [panayotov2015librispeech], using the annotations provided by [tang2023salmonn]. The visual-text case is trained for one epoch, and other cases are trained for three epochs. Rank of low-rank matrices is . The remaining settings are the same as the first stage.
@@
Audio-visual-text: in-domain test set of MUSIC-AVQA and AVE dataset.
Visual-text: public benchmarks: MMEpercep [fu2024mmecomprehensiveevaluationbenchmark], MMBench [liu2024mmbench], POPE [li2023evaluating], SEED-Bench [li2023seedbenchbenchmarkingmultimodalllms].
Speech-text: public benchmarks: MMAUmini-speech [sakshi2024mmau] as well as the foundation subset of
AIR-Benchspeech-en [yang2024air].
Large Language Model. For all three cases, LLaMA-2-7b-Chat [touvron2023llama], LLaMA-3-8B-Instruct [grattafiori2024llama], and Qwen2-7B-Instruct [yang2024qwen2technicalreport] are used as the LLM base model, respectively. For the audio-visual-text case, Qwen2.5-VL-7B-Instruct [bai2025qwen2] is also used as the LLM base model. Throughout the training process, weights of LLM are kept frozen. More experiments of Qwen3 are provided in Appendix A.
@@ 4.1 Audio-visual-text scenario
| LoRAMoE [dou2023loramoe] | 73.57 | 72.81 | |||
| DoRA [liu2024dora] | 73.97 | 72.18 | 1 | 1 | |
| HydraLoRA [tian2024hydralora] | 74.12 | 72.27 | 1 | ||
| Uni-modal LoRA [abouelenin2025phi] | 74.37 | 71.44 | |||
@@
| LoRAMoE [dou2023loramoe] | 73.48 | 72.83 | |||
| DoRA [liu2024dora] | 73.29 | 72.91 | 1 | 1 | |
| HydraLoRA [tian2024hydralora] | 73.14 | 72.59 | 1 | ||
| Uni-modal LoRA [abouelenin2025phi] | 73.62 | 73.14 | |||
@@
| LoRAMoE [dou2023loramoe] | 73.28 | 71.91 | |||
| DoRA [liu2024dora] | 73.37 | 71.06 | 1 | 1 | |
| HydraLoRA [tian2024hydralora] | 73.04 | 71.26 | 1 | ||
| Uni-modal LoRA [abouelenin2025phi] | 73.46 | 72.11 | |||
@@
To validate the effectiveness of our MokA fine-tuning strategy, we compare it with LoRA [hu2022lora] and its variants, including multiple LoRA, LoRAMoE [dou2023loramoe], DoRA [liu2024dora], HydraLoRA [tian2024hydralora], Uni-modal LoRA [abouelenin2025phi]. In addition, we also compare with two additional baselines, whose frameworks are provided in Figure 4. Concretely, the Uni LoRA + MM LoRA strategy employs unimodal low-rank matrices to extract unimodal information independently, while incorporating an additional fully shared multimodal LoRA module to implicitly promote cross-modal interaction. The Uni LoRA + MM LoRA + Gate variant further introduces a gating mechanism to dynamically integrate the outputs of the Uni LoRA and MM LoRA branches for improved fusion. These two baselines incorporate our multimodal-aware basis that ensures both unimodal and cross-modal adaptation, but involve more parameters and offer limited cross-modal interaction. Based on the results in Table 1, we can have the following observations:
Our proposed MokA method achieves the superior overall performance across multiple audio-visual-text datasets, consistently outperforming other baselines and compared methods. While MokA introduces a slight increase in parameter scale compared to standard LoRA, this does not account for the observed performance improvements. Based on the Table 1, multiple LoRA, a baseline that uses A matrices and A matrices, underperforms both standard LoRA and MokA. Simply increasing the number of low-rank matrices does not necessarily lead to better fine-tuning performance. This suggests that MokA’s advantage stems not from parameter quantity, but from its insurance for both unimodal and multimodal adaptation.
@@
| LoRAMoE [dou2023loramoe] | 938.52 | 51.98 | 71.15 | 39.13 | |||
| DoRA [liu2024dora] | 786.47 | 51.31 | 71.07 | 38.96 | 1 | 1 | |
| HydraLoRA [tian2024hydralora] | 774.47 | 47.33 | 70.87 | 38.81 | 1 | ||
| Uni-modal LoRA [abouelenin2025phi] | 992.31 | 51.98 | 72.24 | 39.27 | |||
@@
| LoRAMoE [dou2023loramoe] | 1157.39 | 57.29 | 81.29 | 56.39 | N | N | |
| DoRA [liu2024dora] | 1024.42 | 56.19 | 80.75 | 55.03 | 1 | 1 | |
| HydraLoRA [tian2024hydralora] | 1098.25 | 56.42 | 81.34 | 54.67 | 1 | ||
| Uni-modal LoRA [abouelenin2025phi] | 1189.47 | 57.39 | 81.12 | 56.21 | |||